# Barycenter curvature-dimension condition for extended metric measure spaces

### v1

@@ Abstract
In this survey, we introduce a new curvature-dimension condition for extended metric measure spaces, called Barycenter-Curvature Dimension condition BCD, from the perspective of Wasserstein barycenter.
@@ 1 Optimal transport and displacement convexity
The theory of optimal transport stems from a practical problem:
@@
In the language of mathematics, this problem is to minimize the following quantity:
$I[\pi]=\int_{X\times X}c(x,y)\,{\mathrm{d}}\pi(x,y)~{}~{}~{}\text{for }\pi\in\Pi(\mu,\nu),$
where $c:X\times X\rightarrow[0,\infty]$ is a cost function and $\Pi(\mu,\nu)$ is composed of all probability measure (transport plan) $\pi$ on $X\times X$ whose marginals are $\mu$ and $\nu$. We denote the infimum of $I[\pi]$ by $T_{c}(\mu,\nu)$.
Let $(X,\mathsf{d})$ be a metric space and consider the optimal transport problem with the cost function $c=\mathsf{d}^{2}$. We define the Wasserstein distance $W_{2}$ as $W_{2}^{2}(\mu,\nu):=T_{\mathsf{d}^{2}}(\mu,\nu)$. Evidently, it is an extended metric (allows distance $W_{2}=\infty$) on the space of Borel probability measure $\mathcal{P}(X)$. To make it become a metric, we often restrict the study on subspace $\mathcal{P}_{2}(X,\mathsf{d}):=\left\{\mu\in\mathcal{P}(X):\int_{X}\mathsf{d}^{2}(x,x_{0})\,{\mathrm{d}}x<\infty~{}\text{for some}~{}x_{0}\in X\right\}$. The metric space $(\mathcal{P}_{2}(X,\mathsf{d}),W_{2})$ is called Wasserstein space. The geometry theory of Wasserstein space has a wide range of applications in many fields, including differential equations, geometric and functional inequalities. We refer to [Vil09] for more materials. In this survey, we mainly focus on the connection between optimal transport and the geometry of metric measure spaces.
In his celebrated paper [McC97], McCann studied the convexity of functionals, defined on $\mathcal{P}_{2}(\mathbb{R}^{n})$, with respect to Wasserstein distance. We say that a functional $\mathcal{F}$ is $K$-displacement convex (in the sense of McCann) for some $K\in\mathbb{R}$, if for any $\mu_{0},\mu_{1}\in\mathcal{P}_{2}(\mathbb{R}^{n})$, there is a geodesic ${\mu_{t}}$ in the Wasserstein space, connecting $\mu_{0},\mu_{1}$ such that
@@
There is a family of functionals, called internal energies, which plays crucial role in the study of metric measure spaces. Precisely, the internal energy $\mathcal{H}^{f}_{\mathfrak{m}}$ is defined as
$\mathcal{H}^{f}_{\mathfrak{m}}(\mu):=\left\{\begin{array}[]{ll}\int f(\rho)\,{\mathrm{d}}\mathfrak{m}{}{}{}&\text{if}~{}\mu=\rho\,\mathfrak{m}\\ +\infty&\text{otherwise},\end{array}\right.$
where $f:[0,\infty)\rightarrow[0,\infty]$ is lower semicontinuous, convex and with more than linear growth. When $f=\rho\ln\rho$, $\mathcal{H}^{f}_{\mathfrak{m}}$ is known as Boltzmann entropy, and denoted by ${\rm Ent}_{\mathfrak{m}}$.
Example 1.2 (Boltzmann entropy).
@@
${\rm Ent}_{\mathfrak{m}}(\mu_{t})\leq(1-t){\rm Ent}_{\mathfrak{m}}(\mu_{0})+t{\rm Ent}_{\mathfrak{m}}(\mu_{1}).$
@@ 2 Synthetic theory of Ricci curvature
The link between displacement convexity and geometry of metric measure spaces has been deeply studied in the viewpoint of Ricci curvature. On one hand, Cordero-Erausquin, McCann and Schmuckenschläger [CEMS01] show that the Boltzmann entropy is $K$-displacement convex in the Wasserstein space over a compact Riemannian manifold with Ricci curvature bounded from below by $K$. On the other hand, von Renesse and Sturm [vRS05] prove that displacement convexity actually implies Ricci curvature lower bound. Thus, in the setting of Riemannian manifolds, lower Ricci curvature bounds can be (equivalently) characterized by the displacement convexity of certain functionals. This fact is one of the main motivations and starting points of Lott–Sturm–Villani theory about non-smooth metric measure spaces.
Usually, Ricci curvature appears in three places:
  • A:
    estimate of the Jacobian determinant of the exponential map;
  • B:
    Bochner’s formula;
  • C:
    gradient estimate of heat flow.
For example, Cordero-Erausquin–McCann–Schmuckenschläger and von Renesse–Sturm’s results build an connection between these aspects on Riemannian manifolds. These approaches also play crucial roles in the synthetic theory of Ricci curvature (lower bound), and we will recall some related results in this section.
Three curvature-dimension conditions
A: LSV curvature-dimension condition
A curvature-dimension condition for metric measure spaces, called Lott–Sturm–Villani (LSV) curvature-dimension condition today, is introduced independently by Lott–Villani [LV09] and Sturm [Stu06a]. They adopt Cordero-Erausquin– McCann–Schmuckenschläger’s interpolation inequality [CEMS01] as a definition.
@@
Let $K\in\mathbb{R}$. A geodesic metric measure space $(X,\mathsf{d},\mathfrak{m})$ verifies ${\rm CD}(K,\infty)$ condition, if, for any two $\mu_{0},\mu_{1}\in\mathcal{P}_{2}(X,\mathsf{d})$, there exists a geodesic $\mu_{t}$ such that for all $t\in(0,1)$:
@@
B: BE curvature-dimension condition
Barky and Émery [BÉ85] use Bochner’s inequality as a definition of curvature-dimension bounds, in the setting of Dirichlet form theory.
Definition 2.2.
Let $\mathcal{E}$ be a local and symmetric Dirichlet form on the measure space $(X,\mathfrak{m})$ with dense domain ${\rm D}(\mathcal{E})\in L^{2}(X,\mathfrak{m})$ and $(P_{t})_{t\geq 0}$ be the associated Markov semigroup on $L^{2}(X,\mathfrak{m})$ with generator $\Delta$. On a suitable algebra $\mathcal{A}$ of functions dense in the domain ${\rm D}(\mathcal{E})$, one can define the carré du champ operator $\Gamma$ by
$\Gamma(f,g):=\frac{1}{2}\big{(}\Delta(fg)-f\Delta g-g\Delta f\big{)},~{}~{}~{}f,g\in\mathcal{A}.$
Let $\Gamma_{2}$ be a further iteration of the above procedure:
$\Gamma_{2}(f,g):=\frac{1}{2}\big{(}\Delta\Gamma(f,g)-\Gamma(f,\Delta g)-\Gamma(\Delta f,g)\big{)}.$
We say that the measure space $(X,\mathfrak{m})$ verfies ${\rm BE}(K,\infty)$ condition if the following inequality holds:
$\Gamma_{2}(f,f)\geq K\Gamma(f,f).$ (2.2)
Note that, in case $(X,\mathfrak{m})$ is a Riemannian manifold equipped with its volume measure and $\Delta$ is the Laplace–Beltrami operator, the inequality (2.2) is exactly Bochner’s inequality. We refer to [Led14] and [BGL14] for more discussions on this topic.
C: Evolution Variation Inequality (EVI)
In [DS08], Daneri and Savaré provide a different way to study synthetic curvature-dimension condition in non-compact spaces (see Ambrosio–Erbar–Savaré [AES16] for more discussions in the setting of extended metric measure spaces). They show that a stronger formulation of convexity, called EVI-gradient flow (Evolution Variation Inequality), implies (geodesically) convexity. In particular, one can prove McCann’s displacement convexity by showing the existence of EVI-gradient flow in Wasserstein space.
Definition 2.3 (See [AES16]).
Let $(Y,\mathsf{d}_{Y})$ be an extended metric space and $E:Y\rightarrow\mathbb{R}\cup\left\{+\infty\right\}$ be a lower semi-continuous functional, ${\mathrm{D}(E)}:={\left\{E<+\infty\right\}}$. For any $y_{0}\in\mathrm{D}(E)$, we say that $(0,\infty)\ni t\rightarrow y_{t}\in Y$ is an ${\rm EVI}_{K}$ gradient flow of $E$, starting from $y_{0}$, if it is a locally absolutely continuous curve, such that $y_{t}\overset{\mathsf{d}_{Y}}{\rightarrow}y_{0}$ as $t\rightarrow 0$ and the following inequality is satisfied for any $z\in\mathrm{D}(E)$ satisfying $\mathsf{d}_{Y}(z,y_{t})<\infty$ for some (and then all) $t\in(0,\infty)$:
The $\rm EVI$ gradient flow encodes information not only functional itself but also the structure of the underling space. For example, in a Finsler manifold, the existence of $\rm EVI$ gradient flows of the relative entropy implies that the manifold is Riemannian.
Examples:
The property that the gradient flow of a $K$-convex function $E$ is ${\rm EVI}_{K}$ type, is valid on Euclidean spaces, Hilbert spaces, Alexandrov spaces, ${\rm CAT}$ spaces. For the Wasserstien space $(\mathcal{P}(X),W_{2})$, this property is valid, when $X$ is an Euclidean space, a smooth Riemannian manifold with uniform Ricci lower bound, an Alexandrov space with curvature bounded from below, a Wiener spaces or an $\rm RCD$ space (see §3). We refer the readers to [AGS05, AES16, MS20] for more discussions on this topic. Moreover, we will show its connection with Wasserstein barycenter in Section 3.
In [AGS15], Ambrosio, Gigli and Savaré prove that BE condition, under some natural hypothesis, implies EVI-gradient flow, which then implies a “Riemmanian LSV condition”, the so-called Riemannian curvature-dimension (RCD) condition. Conversely, in [AGS14], they show that $\rm BE$ condition could be deduced from RCD condition. We will discuss this connection in the next subsection.
@@
In [AGS14], Ambrosio, Gigli and Savaré introduce a Riemannian version of the curvature-dimension condition, called Riemannian Curvature-Dimension condition, ruling out Finsler manifolds.
Recall that Cheeger energy (cf. [Che99]) is a functional defined in $L^{2}(X,\mathfrak{m})$ by
@@
Definition 2.4 (${\rm RCD}(K,\infty)$ condition).
Let $K\in\mathbb{R}$. A metric measure space $(X,\mathsf{d},\mathfrak{m})$ verifies Riemannian curvature bounded from below by $K$ (or $(X,\mathsf{d},\mathfrak{m})$ is an ${\rm RCD}(K,\infty)$ space) if it satisfies ${\rm CD}(K,\infty)$ condition and the Cheeger energy ${\rm Ch}$ is a quadratic form (Riemannian structure) in the sense that
${\rm Ch}(f+g)+{\rm Ch}(f-g)=2{\rm Ch}(f)+2{\rm Ch}(g)\qquad\text{for all}~{}f,g\in W^{1,2}(X).$ (2.4)
Formally speaking, an $\rm RCD$ space is a $\rm CD$ space equipped with a Riemannian structure. Remarkably, the above $\rm RCD$ condition can be equivalently defined via gradient flow.
@@
$\int_{X}e^{-c\mathsf{d}(x_{0},x)^{2}}\,{\mathrm{d}}\mathfrak{m}(x)<\infty,\qquad\text{for all }x_{0}\in X\text{ and }c>0,$
and for all $\mu\in\mathcal{P}_{2}(X,\mathsf{d})$ there exists an ${\rm EVI}_{K}$ gradient flow, in $(\mathcal{P}_{2}(X,\mathsf{d}),W_{2})$, of ${\rm Ent}_{\mathfrak{m}}$ starting from $\mu$.
@@ 3 Wasserstein barycenter
Let $\Omega$ be a Borel probability measure on an extended metric space $(Y,\mathsf{d}_{Y})$, a barycenter of $\Omega$ is defined as a minimizer of
@@
and the variance of $\Omega$ is defined as
@@
When $(Y,\mathsf{d}_{Y})$ is a Wasserstein space, a barycenter of $\Omega$ is also called Wasserstein barycenter. This kind of problems, called Wasserstein barycenter problems, draw particular interests in optimal transport and its related fields, as it gives a natural but non-linear way to interpolate between a distribution of measures.
In metric measure geometry setting, Agueh–Carlier [AC11] in Euclidean spaces, Kim–Pass [KP17] and Ma [Ma23] in Riemannian manifolds, Jiang [Jia17] in Alexandrov spaces, establish the existence and uniqueness of Wasserstein barycenter, and the absolutely continuity of the Wasserstein barycenter with respect to the canonical reference measures in these cases. As a consequence, they get the following Wasserstein Jensen’s inequality, which characterizes the convexity of certain functionals via Wasserstein barycenter:
For any $K$-displacement convex functional $\mathcal{F}:\mathcal{P}_{2}(X,\mathsf{d})\to\mathbb{R}\cup\{+\infty\}$ in the sense of McCann [McC97], a probability measure $\Omega$ on $\mathcal{P}_{2}(X,\mathsf{d})$ and its unique barycenter $\bar{\mu}$, it holds
@@
The above inequality can be understood as a general version of “displacement convexity” with respect to Wasserstein barycenter. Motivated by the synthetic theory of Ricci curvature, it is natural to ask:
@@
One may try to imitate the approach used in $\mathbb{R}^{n}$, manifolds and Alexandrov spaces. A barrier lies in the lack of sectional curvature bound, which plays an important role in deriving the regularity of Wasserstein barycenter in their proofs. To overcome this difficulty, we deal with Wasserstein barycenter in a synthetic approach, by making full use of Jensen’s inequality. We observe that the Wasserstein Jensen’s inequality actually encodes the regularity of barycenter. For example, applying $\mathcal{F}={\rm Ent}_{\mathfrak{m}}$ in (3.1), we get
@@
So ${\rm Ent}_{\mathfrak{m}}(\bar{\mu})<+\infty$ in case $\int_{\mathcal{P}_{2}(X,\mathsf{d})}{\rm Ent}_{\mathfrak{m}}(\mu)\,{\mathrm{d}}\Omega(\mu)<+\infty$ and $\Omega$ has finite variance.
Motivated by the work of Daneri and Savaré [DS08], we realize that the existence of $\rm EVI$ gradient flow, introduced in Section 2, implies the Jensen’s inequality.
Theorem 3.1 ($\rm EVI$ implies Jensen’s inequality).
@@
where $y_{t}$ is the ${\rm EVI}_{K}$ gradient flow of E starting from $y$. In particular, if $\mu$ has a barycenter $\bar{y}$ and there is an ${\rm EVI}_{K}$ gradient flow of $E$ starting from $\bar{y}$, then we have Jensen’s inequality:
@@
The following proposition, proved in [DS08, Proposition 3.1] (see also [AES16]), provides an integral version of ${\rm EVI}$ type gradient flows. We will use this formulation to prove Jensen’s inequality.
Proposition 3.2 (Integral version of $\rm EVI$).
Let $E$, $K$ and $(y_{t})_{t>0}$ be as in Definition 2.3, then $(y_{t})_{t>0}$ is an ${\rm EVI}_{K}$ gradient flow if and only if it satisfies
$\frac{e^{K(t-s)}}{2}\mathsf{d}_{Y}^{2}(y_{t},z)-\frac{1}{2}\mathsf{d}_{Y}^{2}(y_{s},z)\leq I_{K}(t-s)\big{(}E(z)-E(y_{t})\big{)},\quad\forall 0\leq s\leq t,$ (3.5)
@@
Combining Theorem 2.5 and Theorem 3.1 we establish the Wasserstein Jensen’s inequality in $\rm RCD$ spaces.
Example 3.3.
@@
Furthermore, the geometry of some extended measure spaces, such as Wiener spaces [AES16] and configuration spaces [EH15], can also be characterized via $\rm EVI$ gradient flows with initial data in ${\rm D}({\rm Ent})$. Concerning the application of Theorem 3.1 for extended metric measure spaces, it is worth mentioning that one should take care of the existence of $\rm EVI$ gradient flows. In general, this is a highly non-trivial problem.
Example 3.4.

### v2

@@ Abstract
In this survey, we introduce a new curvature-dimension condition for extended metric measure spaces, called Barycenter-Curvature Dimension condition, from the perspective of Wasserstein barycenter.
@@ 1 Optimal transport and displacement convexity
Optimal transport theory stems from a practical problem:
@@
In mathematics, this can be stated as a minimization problem:
$T_{c}(\mu,\nu)=\min_{\pi\in\Pi(\mu,\nu)}\int_{X\times X}c(x,y)\,{\mathrm{d}}\pi(x,y)$
where $c:X\times X\rightarrow[0,\infty]$ is a cost function and $\Pi(\mu,\nu)$ is composed of all probability measure $\pi$ on $X\times X$ whose marginals are $\mu$ and $\nu$.
Let $(X,\mathsf{d})$ be a metric space and consider the optimal transport problem with the cost function $c=\mathsf{d}^{2}$. The Wasserstein distance $W_{2}$ is defined by $W_{2}^{2}(\mu,\nu):=T_{\mathsf{d}^{2}}(\mu,\nu)$. It is immediate to see that $W_{2}$ is an extended metric (we allow $W_{2}(\mu,\nu)=\infty$) on the space of Borel probability measures $\mathcal{P}(X)$. To make it become a metric, we often restrict the study on a subspace
$\mathcal{P}_{2}(X,\mathsf{d}):=\left\{\mu\in\mathcal{P}(X):\int_{X}\mathsf{d}^{2}(x,x_{0})\,{\mathrm{d}}x<\infty~{}\text{for some}~{}x_{0}\in X\right\}.$
The metric space $(\mathcal{P}_{2}(X,\mathsf{d}),W_{2})$ is called Wasserstein space. The geometry theory of Wasserstein space has a wide range of applications in many fields, including differential equations, geometric and functional inequalities. We refer to [AG11] and [Vil09] for a comprehensive discussion about optimal transport and Wasserstein space. In this survey, we mainly focus on the connection between optimal transport and the geometry of metric measure spaces.
In his celebrated paper [McC97], McCann studied the convexity of functionals on $(\mathcal{P}_{2}(\mathbb{R}^{n},|\cdot|),W_{2})$. We say that a functional $\mathcal{F}$ is $K$-displacement convex (in the sense of McCann) for some $K\in\mathbb{R}$, if for any $\mu_{0},\mu_{1}\in\mathcal{P}_{2}(\mathbb{R}^{n},|\cdot|)$, there exists a geodesic $\{\mu_{t}\}$ in the Wasserstein space, connecting $\mu_{0}$ and $\mu_{1}$, such that
@@
There is a family of functionals, called internal energies, which plays crucial role in the study of metric measure spaces. In physics, this kind of functionals are more known as entropies since they have a thermodynamic interpretation of internal energy of a gas. Precisely, the internal energy $\mathcal{H}^{U}_{\mathfrak{m}}$ is defined as
$\mathcal{H}^{U}_{\mathfrak{m}}(\mu):=\left\{\begin{array}[]{ll}\int U(\rho)\,{\mathrm{d}}\mathfrak{m}{}{}{}&\text{if}~{}\mu=\rho\,\mathfrak{m}\\ +\infty&\text{otherwise},\end{array}\right.$
where $U:[0,\infty)\rightarrow[0,\infty]$ is lower semicontinuous, convex and with super-linear growth. When $U(x)=x\ln x$, $\mathcal{H}^{U}_{\mathfrak{m}}$ is known as Boltzmann entropy, and it is usually denoted by ${\rm Ent}_{\mathfrak{m}}$. We refers the readers to [AGS08] and [San15] for more functionals and applications.
Proposition 1.2 (Boltzmann entropy, Theorem 2.2 in [McC97]).
@@
${\rm Ent}_{\mathfrak{m}}(\mu_{t})\leq(1-t){\rm Ent}_{\mathfrak{m}}(\mu_{0})+t{\rm Ent}_{\mathfrak{m}}(\mu_{1})~{}~{}~{}t\in[0,1]$
for any geodesic $\{\mu_{t}\}$ in the Wasserstein space.
@@ 2 Synthetic theory of Ricci curvature
Since the seminal paper of Otto and Villani [OV00], the link between displacement convexity and geometry of metric measure spaces has been deeply studied from the perspective of Ricci curvature. On one hand, Cordero-Erausquin, McCann and Schmuckenschläger [CEMS01] showed that the Boltzmann entropy is $K$-displacement convex in the Wasserstein space over a compact Riemannian manifold with Ricci curvature bounded from below by $K$. On the other hand, von Renesse and Sturm [vRS05] proved that displacement convexity actually implies lower Ricci curvature bounds. Thus, in the setting of Riemannian manifolds, lower Ricci curvature bounds can be characterized by the displacement convexity of certain functionals. This fact is one of main motivation and starting points of Lott–Sturm–Villani theory about non-smooth metric measure spaces.
Curvature-dimension condition
A synthetic curvature-dimension condition for metric measure spaces, called CD condition today, was introduced independently by Lott–Villani [LV09] and Sturm [Stu06a]. They adopt the interpolation inequality of Cordero-Erausquin, McCann and Schmuckenschläger [CEMS01] as a definition.
@@
Let $K\in\mathbb{R}$. We say that a metric measure space $(X,\mathsf{d},\mathfrak{m})$ verifies ${\rm CD}(K,\infty)$ condition, if for any $\mu_{0},\mu_{1}\in\mathcal{P}_{2}(X,\mathsf{d})$, there exists a geodesic $\mu_{t}$ such that for all $t\in(0,1)$:
@@
It is worth to mention that ${\rm CD}(K,\infty)$ is a dimensionless curvature condition. There is a variant, called ${\rm CD}(K,N)$ condition, which has a dimensional parameter $N$. We refer to [Stu06b] and [Vil09] for the definition.
@@
In [Oht09], Ohta proved that some Finsler manifolds satisfy $\rm CD$ condition. Later, Ambrosio, Gigli and Savaré, in [AGS14], introduced a Riemannian variant of the curvature-dimension condition, called Riemannian Curvature-Dimension condition ($\rm RCD$ for short), ruling out Finsler manifolds.
Recall that Cheeger energy (cf. [Che99]) is a functional defined on $L^{2}(X,\mathfrak{m})$:
@@
Note that, a smooth Finsler manifold is Riemannian if and only if its Sobolev space $W^{1,2}$ is a Hilbert space; see [Gig23, Section 4]. This fact motivates the following definition, proposed by Gigli in [Gig15].
Definition 2.2 (Infinitesimally Hilbertian space).
We say that a metric measure space $(X,\mathsf{d},\mathfrak{m})$ is an infinitesimally Hilbertian space if $W^{1,2}(X,\mathsf{d},\mathfrak{m})$ is Hilbert. Equivalently, the Cheeger energy ${\rm Ch}$ is a quadratic form in the sense that
${\rm Ch}(f+g)+{\rm Ch}(f-g)=2{\rm Ch}(f)+2{\rm Ch}(g)\qquad\text{for all}~{}f,g\in W^{1,2}(X,\mathsf{d},\mathfrak{m}).$
We can then give the following definition:
Definition 2.3 (${\rm RCD}(K,\infty)$ condition, see [AGS14] and [AGMR15]).
Let $K\in\mathbb{R}$. A metric measure space $(X,\mathsf{d},\mathfrak{m})$ verifies Riemannian curvature bounded from below by $K$ (or $(X,\mathsf{d},\mathfrak{m})$ is an ${\rm RCD}(K,\infty)$ space) if it satisfies ${\rm CD}(K,\infty)$ condition and $W^{1,2}(X,\mathsf{d},\mathfrak{m})$ is a Hilbert space.
Evolution Variation Inequality (EVI)
In [DS08], Daneri and Savaré provided an alternative way to study synthetic curvature-dimension condition in possibly infinite dimensional, non-compact spaces (see Ambrosio, Erbar and Savaré’s paper [AES16] for more discussion in the setting of extended metric measure spaces). They established the existence of a stronger formulation of gradient flows, called ${\rm EVI}_{K}$ (Evolution Variation Inequality) gradient flows, on Riemannian manifolds.
Definition 2.4 (See [AES16]).
Let $(Y,\mathsf{d}_{Y})$ be an extended metric space and $E:Y\rightarrow\mathbb{R}\cup\left\{+\infty\right\}$ be a lower semi-continuous functional. Denote ${\mathrm{D}(E)}:={\left\{E<+\infty\right\}}$. For any $y_{0}\in\mathrm{D}(E)$, we say that $(0,\infty)\ni t\rightarrow y_{t}\in Y$ is an ${\rm EVI}_{K}$-gradient flow of $E$, starting from $y_{0}$, if it is a locally absolutely continuous curve, such that $y_{t}\overset{\mathsf{d}_{Y}}{\rightarrow}y_{0}$ as $t\rightarrow 0$ and the following inequality is satisfied: for any $z\in\mathrm{D}(E)$ satisfying $\mathsf{d}_{Y}(z,y_{t})<\infty$ for some (and then all) $t\in(0,\infty)$:
It has been proved in [DS08] that this stronger formulation of gradient flows implies (geodesic) convexity. As an application, one can prove McCann’s displacement convexity by showing the existence of ${\rm EVI}_{K}$-gradient flows in Wasserstein space.
The ${\rm EVI}_{K}$-gradient flow encodes information not only about the functional $E$ itself but also the structure of the underling space. On one hand, Ohta and Sturm [OS09] proved that the heat flow on a Finsler manifold is linear if and only if it is Riemannian. On the other hand, by Daneri and Savaré [DS08], we know the existence of ${\rm EVI}_{K}$-gradient flows of the relative entropy implies the linearity of the heat flow. This suggests that ${\rm EVI}_{K}$-gradient flow is closely related to the Riemannian (infinitesimally Hilbertian) structure of the underlying space. Remarkably, the $\rm RCD$ condition can be equivalently defined via heat flow.
@@
$\int_{X}e^{-c\mathsf{d}(x_{0},x)^{2}}\,{\mathrm{d}}\mathfrak{m}(x)<\infty,\qquad\text{for some }x_{0}\in X\text{ and }c>0,$
and for all $\mu\in\mathcal{P}_{2}(X,\mathsf{d})$ there exists an ${\rm EVI}_{K}$-gradient flow of ${\rm Ent}_{\mathfrak{m}}$, in $(\mathcal{P}_{2}(X,\mathsf{d}),W_{2})$, starting from $\mu$.
Remark 2.6.
There is another well known synthetic curvature-dimension condition, called Bakry-Émery condition (BE condition for short), named after Bakry and Émery [BE85] (see [Led14, BGL14] for more discussion on this topic). In [AGS15], Ambrosio, Gigli and Savaré proved that BE condition, under some natural hypothesis, implies the existence of ${\rm EVI}_{K}$-gradient flows, which then implies ${\rm RCD}$ condition. Conversely, in [AGS14], they show that $\rm BE$ condition could be deduced from RCD condition.
Example 2.7.
The existence of the ${\rm EVI}_{K}$-gradient flow of a $K$-convex function $E$, is valid on Euclidean spaces, Hilbert spaces, Alexandrov spaces, ${\rm CAT}$ spaces; see [AGS08] and [MS20, Section 3.4].
Example 2.8.
For the Wasserstien space $(\mathcal{P}_{2}(X,\mathsf{d}),W_{2})$ over metric measure spaces, the existence of the ${\rm EVI}_{K}$-gradient flow is valid, when $X$ is an Euclidean space, a smooth Riemannian manifold with uniform Ricci lower bound, an Alexandrov space with curvature bounded from below or an $\rm RCD$ space; see also [AGS14] and [GKO13].
Example 2.9.
The existence of ${\rm EVI}_{K}$-gradient flows also holds on the Wasserstien space $(\mathcal{P}(X),W_{2})$ (we denote the Wasserstein space over an extended metric space by $(\mathcal{P}(X),W_{2})$ since $\mathcal{P}_{2}(X,{\mathrm{d}})$ makes no sense) over extended metric measure spaces, such as a Wiener space [FSS10] and configuration spaces [EH15], see [AES16, Section 13] for more discussion. We will discuss its connection with Wasserstein barycenter problem in Section 3.
@@ 3 Wasserstein barycenter
We say that $(Y,\mathsf{d}_{Y})$ is an extended metric space if $\mathsf{d}_{Y}:Y\times Y\to[0,+\infty]$ is a symmetric function satisfying the triangle inequality, with $\mathsf{d}_{Y}(x,y)=0$ if and only if $x=y$. Let $\Omega$ be a Borel probability measure on $(Y,\mathsf{d}_{Y})$. A minimizer of the function
@@
is called a barycenter of $\Omega$, and the variance of $\Omega$ is defined as
@@
In the Wasserstein space $(\mathcal{P}_{2}(X,\mathsf{d}),W_{2})$, a barycenter of $\Omega$ is also called Wasserstein barycenter. This kind of problems, called Wasserstein barycenter problems, draw particular interests in optimal transport and its related fields, as it gives a natural way to average several measures. We remark that the barycenter problem is independent of the linearity of the underling space, and is essentially a non-linear problem.
In Euclidean space, Agueh–Carlier [AC11] established the well-posedness (i.e. existence, uniqueness and the absolute continuity) of the Wasserstein barycenter problem and prove the following Jensen’s inequality.
Theorem 3.1 ([AC11]).
In $(\mathbb{R}^{n},|\cdot|,\mathcal{L}^{n})$, the Boltzmann entropy satisfies the Wasserstein Jensen’s inequality:
${\rm Ent}_{\mathcal{L}^{n}}(\bar{\mu})\leq\int_{\mathcal{P}_{2}(\mathbb{R}^{n},|\cdot|)}{\rm Ent}_{\mathcal{L}^{n}}(\mu)\,{\mathrm{d}}\Omega(\mu)$ (3.1)
where $\Omega$ is supported in finitely many measures over $\mathcal{P}_{2}(\mathbb{R}^{n},|\cdot|)$ with finite variance and $\bar{\mu}$ is a barycenter of $\Omega$.
Moreover, Kim–Pass [KP17] extended the above results to compact Riemannian manifolds; Ma [Ma23] improves Kim–Pass’s results in Riemannian manifolds when $\Omega$ is general distributions (with infinite support); Jiang [Jia17] proved the well-posedness in Alexandrov spaces. As a consequence, the following generalized Wasserstein Jensen’s inequality has been proved, which characterizes the convexity of certain functionals via Wasserstein barycenter.
Let $\mathcal{F}:\mathcal{P}_{2}(X,\mathsf{d})\to\mathbb{R}\cup\{+\infty\}$ be a $K$-displacement convex functional in the sense of McCann [McC97], $\Omega$ be a probability measure on $\mathcal{P}_{2}(X,\mathsf{d})$. We say that Wasserstein Jensen’s inequality holds if there is a barycenter $\bar{\mu}$ of $\Omega$ such that
@@
The above inequality can be understood as a generalization of “displacement convexity”, from the perspective of Wasserstein barycenter. Motivated by the synthetic theory of curvature-dimension condition, it is natural to ask:
@@
One may try to mimic the argument used for $\mathbb{R}^{n}$ or Riemannian manifolds: firstly prove the regularity of Wasserstein barycenter, then prove the Jensen’s inequality using the first order balance condition. A huge barrier lies in the lack of sectional curvature bound, which plays a crucial role in deriving the regularity of Wasserstein barycenter. A new feature in our work is that we deal with Wasserstein barycenter in a synthetic approach, by treating Jensen’s inequality as an “a priori estimate”, so that we do not need to prove the regularity of Wasserstein barycenter at first. In fact, the Wasserstein Jensen’s inequality already encodes the regularity of barycenter. To see this, applying $\mathcal{F}={\rm Ent}_{\mathfrak{m}}$ in (WJI), we get
@@
So ${\rm Ent}_{\mathfrak{m}}(\bar{\mu})<+\infty$ in case that $\int_{\mathcal{P}_{2}(X,\mathsf{d})}{\rm Ent}_{\mathfrak{m}}(\mu)\,{\mathrm{d}}\Omega(\mu)<+\infty$ and $\Omega$ has finite variance.
Motivated by Daneri and Savaré [DS08], we realize that the existence of ${\rm EVI}_{K}$ gradient flow implies the Jensen’s inequality.
Theorem 3.2 (${\rm EVI}_{K}$ implies Jensen’s inequality).
@@
where $\{y_{t}\}$ is the ${\rm EVI}_{K}$ gradient flow of E starting from $y$ and $I_{K}(t)$:=$\int_{0}^{t}e^{Kr}\,{\mathrm{d}}r$. In particular, if $\mu$ has a barycenter $\bar{y}$ and there is an ${\rm EVI}_{K}$ gradient flow of $E$ starting from $\bar{y}$, then we have the Jensen’s inequality:
@@
The following proposition, proved in [DS08, Proposition 3.1] (see also [AES16]), provides an integral version of ${\rm EVI}_{K}$ type gradient flows. We will use this formulation to prove Jensen’s inequality.
Proposition 3.3 (Integral version of ${\rm EVI}_{K}$).
Let $E$, $K$ and $\{y_{t}\}_{t>0}$ be as in Definition 2.4, then $\{y_{t}\}_{t>0}$ is an ${\rm EVI}_{K}$ gradient flow if and only if it satisfies
$\frac{e^{K(t-s)}}{2}\mathsf{d}_{Y}^{2}(y_{t},z)-\frac{1}{2}\mathsf{d}_{Y}^{2}(y_{s},z)\leq I_{K}(t-s)\big{(}E(z)-E(y_{t})\big{)},\quad\forall~{}0\leq s\leq t,$ (3.6)
@@
Combining Theorem 2.5 and Theorem 3.2 we establish Wasserstein Jensen’s inequality in $\rm RCD$ spaces.
Corollary 3.4.
@@
Furthermore, on some extended metric measure spaces, such as abstract Wiener spaces [AES16] and configuration spaces [EH15], ${\rm EVI}_{K}$ gradient flow of the relative entropy is well-posed in the sense of Definition 2.4. To apply Theorem 3.2 for extended metric measure spaces, one should take care of the existence of the gradient flow from a point with finite distance to the domain of the relative entropy. In general, this is a highly non-trivial problem.
Corollary 3.5 (See [FU04] and [FSS10]).